# Active Learning-Based Optimization of Hydroelectric Turbine Startup to Minimize Fatigue Damage

### v1

@@ 4.2 Results
The sequential startup parameters $\theta^{j}$ are given in Table 2, together with the largest observed strain cycle. Figure 9 provides the trajectories projected on the $(o,\omega)$ space. The color map shows the latest trained virtual sensor’s predictions of $\hat{\mu}+\hat{\sigma}$ for the upper (Figure 9(a)) and lower (Figure 9(b)) bounds of the strain envelope. Figure 10 shows the optimization process’s measured strains on the various startups over time.
@@
The active learning trajectories already show a significant strain reduction. Clearly, they follow the lowest opening rate of Initial 2, and catch up for time afterwards. Interestingly, the startup duration $t_{\mathrm{st}}$ is always shorter than the time constraint. This could be due to the small penalty added to time as seen in Equation (19). The Optimal 1 startup is even better, leading to a 42% reduction compared to the best standard startup. Finally, Optimal 2 contends with the need to open the vanes more widely to reach the 60 s time constraint. It still attempts to keep $r_{o}$ lower than most initial trajectories, preferring instead to increase $o_{\mathrm{init}}$ to catch up in time. Optimal 2 still achieves a 26% reduction in strain amplitude compared to the standard parameters, with equal time to reach synchronous speed.
(a) Strain envelope upper bound. White (close to 0) is better.
(b) Strain envelope lower bound. White (close to 0) is better.
Figure 9: Real turbine dynamic trajectories mapped on the opening - rotational speed space. The color map represents the latest trained virtual sensor’s predictions of $\hat{\mu}+\hat{\sigma}$ for the upper and lower bounds of the strain envelope.
@@ 5 Discussion
The results show that the proposed active learning and optimization method is promising and could be applied systematically as part of the commissioning of a new turbine. However, several questions remain to be studied. First, the relevance of the active learning phase depends on the optimization landscape and the number of initial trajectories. In the presented use case, the landscape was smooth and the initial trajectories covered a good portion of the space. The epistemic uncertainty values were low, and the active learning phase, including the results, was not significantly different from the optimization phase. Second, our approach was applied to the strain amplitude measured at a single strain gauge. It is possible that minimizing the strain at one location could increase the strain at another location. A sounder approach would perform optimization over all the installed gauges. The challenges in achieving such a generalized optimization are the computation requirements and robustness to potential defects occurring in strain gauges during the optimization campaign. Third, the water head may influence the turbine strain during startup. During a measurement campaign taking place over a few days, this condition can remain quite constant, but this is not the case during turbine operation. We hypothesize that parameters leading to lower strains for a given water head generalize to other conditions. which remains to be confirmed.
@@ Appendix A Notation table
$T_{\mathrm{st}}$ Maximum time to reach $\omega_{S}$
@@
$n_{\{D,M,e,S\}}$ time steps in trajectory
@@
$\hat{\mu}^{\{u,l\}}$ Mean prediction
$\hat{\sigma}^{\{u,l\}}$ Aleatoric standard deviation prediction
@@
$c(\theta)$ Cost function

### v2

@@ 4.2 Results
The sequential startup parameters $\theta^{j}$ are given in Table 2, together with the largest observed strain cycle. Figure 10 provides the trajectories projected on the $(o,\omega)$ space. The color map shows the latest trained virtual sensor’s predictions of $\hat{\mu}+\hat{\sigma}$ for the upper (Figure 10a) and lower (Figure 10b) bounds of the strain envelope. The highest amplitudes of the strain envelope are observed in operating zone with low rotational speed and higher vane opening, similar to the results presented by Gagnon et al. [24]. This behavior may be explained by the increased stochasticity of the flow in this operating regime, originating from the shedding and collapse of vortices as proposed by Morissette and Nicolle [9]. However, a more comprehensive understanding of this correlation would require an in-depth flow analysis through unsteady CFD-FEA of different startup scenarios, which is beyond the scope of the present paper.
Figure 11 shows the optimization process’s measured strains on the various startups over time.
@@
The active learning trajectories already show a significant strain reduction. Clearly, they follow the lowest opening rate of Initial 2, and catch up for time afterwards. Interestingly, the startup duration $t_{\mathrm{st}}$ is always shorter than the time constraint. This could be due to the small penalty added to time as seen in Equation (19). The Optimal 1 startup is even better, leading to a 42% reduction compared to the best standard startup. Finally, Optimal 2 contends with the need to open the vanes more widely to reach the 60 s time constraint. It still attempts to keep $r_{o}$ lower than most initial trajectories, preferring instead to increase $o_{\mathrm{init}}$ to catch up in time. Optimal 2 still achieves a 26% reduction in strain amplitude compared to the standard parameters, with equal time to reach synchronous speed. The strains measured during the optimal trajectories are shown in higher temporal resolution in figure 12.
Figure 10: Measured turbine dynamic trajectories mapped on the opening - rotational speed space. The color map (white is better) represents the latest trained virtual sensor’s predictions of $\hat{\mu}+\hat{\sigma}$ for the upper (a) and lower bounds (b) of the strain envelope.
Figure 12: Comparison of measured strain envelopes for initial (Ini 5) and optimized startups (Opt 1 is limited to $\bar{t}_{\mathrm{st}}=90$s, Opt 2 is limited to $\bar{t}_{\mathrm{st}}=60$s). While the optimization process is also able to reduce the strain amplitude by 26% for similar durations to the initial startup (Opt 2), it is also able to take advantage of a longer startup time to improve the performance to 42% (Opt 1).
@@ 5 Discussion
The results show that the proposed active learning and optimization method is promising and could be applied systematically as part of the commissioning of a new turbine. An interesting aspect of the method is that the learned model, which can be subject to accuracy errors and is usually delicate to implement in critical industrial settings, is only used to guide the optimization process. The actual strain is measured experimentally on the turbine, guaranteeing the performance of the optimized parameters. However, several questions remain to be studied. First, the relevance of the active learning phase depends on the optimization landscape and the number of initial trajectories. In the presented use case, the landscape was smooth and the initial trajectories covered a good portion of the space. The epistemic uncertainty values were low, and the active learning phase, including the results, was not significantly different from the optimization phase. Second, our approach was applied to the strain amplitude measured at a single strain gauge. It is possible that minimizing the strain at one location could increase the strain at another location. A sounder approach would perform optimization over all the installed gauges. The challenges in achieving such a generalized optimization are the computation requirements and robustness to potential defects occurring in strain gauges during the optimization campaign. Third, the water head may influence the turbine strain during startup. During a measurement campaign taking place over a few days, this condition can remain quite constant, but this is not the case during turbine operation. We hypothesize that parameters leading to lower strains for a given water head generalize to other conditions, which remains to be confirmed.
@@ Appendix A Notation table
$\bar{t}_{\mathrm{st}}$ Maximum time to reach $\omega_{S}$
@@
$T_{turb}$ Turbine torque
$J$ Inertia of rotating components
$q_{4\cdots N_{eq}}$ Internal speed governor model variables
@@
$n_{\{D,M,e,S\}}$ # time steps in trajectory
@@
[…]
@@
$\mathcal{L}$ Objective function for strain
$c(\theta)$ Cost function for the optimizer